# Demystifying Spatial Confounding

### v2

@@ 3.1 Bias in the spatial model
Thus, the bias has the intuitive interpretation as the correlation between $\mathbf{x}$ and $\mathbf{z}$ times the size of $\mathbf{z}$ relative to the size of $\mathbf{x}$ in the metric defined by $\bm{\mathbf{\Sigma}}^{-1}$. That is, in this metric, the bias gets large exactly when the covariate and the confounder are very correlated and the confounder is relatively large compared to the covariate. The sign of the bias is the same as the sign of the correlation, i.e. the sign of $\langle\mathbf{x},\mathbf{z}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}$. The proof of the proposition is in the appendix.
Combining Proposition 1 with the results of Section 2 we get a better understanding of how the bias behaves. Let $\bm{\xi}^{x}=(\bm{\xi}^{xT}_{\mathrm{ns}},\bm{\xi}^{xT}_{\mathrm{sp}})^{T}$ denote the coordinates of $\mathbf{x}$ in the eigenbasis $\mathbf{U}$ such that $\mathbf{x}=\mathbf{U}\bm{\xi}^{x}$, i.e. a non-zero entry in $\bm{\xi}^{x}$ means that the corresponding basis vector is included in $\mathbf{x}$. Similarly, let $\mathbf{z}=\mathbf{U}\bm{\xi}^{z}$. Since $\mathbf{z}$ is spatial, $\bm{\xi}^{z}_{\mathrm{ns}}=\mathbf{0}$. It is now straightforward to prove the following result (see the Web Appendix for the proof).
Corollary 1.
@@
The expression for the bias is not symmetric in $\mathbf{x}$ and $\mathbf{z}$. In particular, it is not affected by components in $\mathbf{z}$, i.e. with $\xi_{\mathrm{sp},i}^{z}\neq 0$, for which the corresponding component in $\mathbf{x}$ is 0. Thus, any unmeasured spatial variation in the response variable that is independent of the covariate can be ignored for the purposes of the bias. However, such components may affect the overall fit of the model and the estimates of the parameters $\lambda$ and $\sigma^{2}$ which affect the weights $w_{i}$.
@@ 3.2 The role of spatial frequencies
Due to the differences in the weights $w_{i}$, different spatial frequencies affect the bias differently. Looking at the numerator in Corollary 1, the larger the weight, the larger the contribution to the bias. So confounding at high frequencies (where weights are close to $1$) induce larger bias, while confounding at low frequencies may give negligible contribution to the bias. Confounding at unpenalised spatial frequencies does not lead to bias (as they have weights 00), and non-spatial components in $\mathbf{x}$ do not contribute to the numerator either. For all confounded frequencies in $\mathbf{x}$ with $w_{i}\neq 0$, the contribution to the bias is proportional to the size $\xi_{\mathrm{sp},i}^{z}$ of the confounder at that frequency. Contributions to the bias from different eigenvectors can partially cancel each other out as they can have different signs.
@@ 3.3 Comparison to the non-spatial model
Irrespective of the size of bias, the non-spatial model (4) is not an appropriate analysis model for the data generation process (2), as the residual spatial variation in the response data violates the assumptions of the model. Indeed, this is the reason for including spatial random effects. Nevertheless, we consider how the bias in model (4) compares to the spatial model (3) as this has been a focus of spatial confounding research in the past. The following corollary is proved in the appendix.
Corollary 2.
@@
Thus, the bias in the non-spatial model differs to that of the spatial model only in that all spatial frequencies have weights $1$ rather than $w_{i}$. Broadly speaking, bias in the non-spatial model reflects the overall correlation between $\mathbf{x}$ and $\mathbf{z}$, whereas bias in the spatial model depends mostly on the correlation at high frequencies. However, as the weights affect both the numerator and denominator of the bias, in general, the relationship between the two models is not straightforward. A detailed computation of the difference and how it is affected by different types of behaviours is given in Web Appendix 1.4. Although the bias tends to be larger in the non-spatial model than in the spatial model, there are scenarios where it is the other way around.
In the past, the estimate in model (4) was often mistakenly assumed to be unbiased and, therefore, a large difference between this and the estimate in (3) was seen as a sign of confounding bias in the spatial model (and, conversely, a small difference a sign of no confounding). However, Corollary 2 shows that a difference between these estimates tends to arise when the covariate is dominated by low frequencies (as these have weights most different from 1) and is not directly related to the existence or otherwise of confounding bias.
@@ 3.4 Dependence on the smoothing parameter
As bias in the spatial model is caused by smoothing, intuitively, we would expect larger values of the parameter $\lambda$ to lead to more bias when all other components are held fixed. However, as $\lambda$ affects both the numerator and the denominator of the expression in Corollary 1, there can be situations where this is not the case. In Web Appendix 1.5 we provide a detailed analysis which shows that, although the behaviour of the bias varies, when there is sufficient non-spatial or unconfounded high frequency components in $\mathbf{x}$, the size of the bias broadly increases with $\lambda$ (in line with intuition). The following corollary is also proved in the Web Appendix.
Corollary 3.
@@
Thus, in the limit $\lambda\rightarrow\infty$, i.e. an entirely smoothed spatial effect, the bias in the spatial model agrees with that of the non-spatial model, except that any contributions from unpenalised spatial components of $\mathbf{x}$ are 00.
In the limit $\lambda\rightarrow 0$, i.e. without smoothing, the behaviour of the bias depends on whether or not the covariate has non-spatial information. If non-spatial information is present, the bias can be eliminated by setting $\lambda=0$, though this is usually undesirable as it is likely lead to an inferior model fit. Similarly, if the proportion of unconfounded high frequency components in $\mathbf{x}$ is large, then the size of the bias reduces as $\lambda\rightarrow 0$ and the limit is close to 00. But in the absence of non-spatial or unconfounded high frequency information, the limit could become large and even exceed the bias of the non-spatial model.
@@ 4 Simulation studies
where $\beta\in\mathbb{R}$, $\bm{\epsilon}^{y}\sim N(\mathbf{0},\sigma^{2}\mathbf{I})$, $\bm{\epsilon}^{x}\sim N(\mathbf{0},\sigma^{2}_{x}\mathbf{I})$ for $\sigma,\sigma_{x}>0$. In this specification, $\bm{\epsilon}^{x}$ represents the non-spatial component of $\mathbf{x}$, while $\mathbf{z}$ and $\mathbf{z}^{x}$ are constructed as linear combinations of two spatial fields $\mathbf{z}_{\mathrm{sp,l}}$ and $\mathbf{z}_{\mathrm{sp,h}}$, which are low and high frequency, respectively, i.e. $\mathbf{z}=\xi_{\mathrm{sp,l}}^{z}\mathbf{z}_{\mathrm{sp,l}}+\xi_{\mathrm{sp,h}}^{z}\mathbf{z}_{\mathrm{sp,h}}$ and $\mathbf{z}^{x}=\xi_{\mathrm{sp,l}}^{x}\mathbf{z}_{\mathrm{sp,l}}+\xi_{\mathrm{sp,h}}^{x}\mathbf{z}_{\mathrm{sp,h}}$ for $\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z},\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x}\in\mathbb{R}$. Unless stated otherwise, we let $\beta=0.5$, $\sigma_{x}=\sigma=1$.
For the regressions in this section, we represent the spatial effect as a thin plate regression spline (TPRS), a computationally efficient reduced rank version of a thin plate spline (Wood,, 2003). The data generating process matches the analysing spatial model which is therefore correctly specified. (In Web Appendix 4, we repeat the simulation where the data generating process is a Gaussian process, so in addition to the spatial confounding bias, there may be further mis-specification bias.) In order to generate $\mathbf{z}_{\mathrm{sp,l}}$ and $\mathbf{z}_{\mathrm{sp,h}}$, we start by considering a mean-zero Gaussian process $\bm{\gamma}$ with exponential covariance structure following $C(h)=\exp(-h/\kappa)$ such data $h=\lVert\mathbf{s}-\mathbf{s}^{\prime}\rVert$ for $\mathbf{s},\mathbf{s}^{\prime}\in[0,1]\times[0,1]$, which is assumed for simplicity to have variance of 1. We set $\kappa=0.1$ corresponding to a spatial range of approximately 0.3. We fit thin plate regression splines to the generated $\bm{\gamma}$ with 10 and 800 basis functions, respectively, and use the corresponding fitted values as $\mathbf{z}_{\mathrm{sp,l}}$ and $\mathbf{z}_{\mathrm{sp,h}}$. Thin plate regression spline models are implemented in the R-package mgcv. We use this implementation with generalised cross-validation as the smoothness selection criterion to compare the results of models fitted to simulated data for which we know the true underlying covariate and spatial dependence. We consider 600 basis functions for the thin plate regression splines in the data analysing model.
@@ 4.1 Scenario 1: Confounding at high frequencies
As expected, Figure 1 (top left) shows significant bias in both models under this scenario, increasing linearly in size with $\xi_{\mathrm{sp,h}}^{z}$. The mean squared error (MSE) of fitted values compared to the true expectation of $\mathbf{y}$ is shown in Web Appendix 3.1 and is always lower for the spatial model, indicating a better fit, which is not surprising as the spatial model is correctly specified, whereas the non-spatial model is mis-specified.
Figure 1: Bias of $\widehat{\beta}$ and $\widehat{\beta}_{\mathrm{ns}}$ (top) in the spatial and non-spatial (ns) models under Scenario 1 (top, left), Scenario 2 (top, right), Scenario 3 (bottom, left), Scenario 4 (bottom, right). The true value of $\beta$ is 0.5.
@@ 4.2 Scenario 2: Confounding at low frequencies
We reconsider Scenario 1 but with high and low frequencies swapped, i.e. $(\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x})=(1,1)$, $\xi_{\mathrm{sp,h}}^{z}=0$ and $\xi_{\mathrm{sp,l}}^{z}\in\{0.2,0.5,1\}$. Since confounding is now at low frequencies, i.e. $\xi_{\mathrm{sp,l}}^{x}\xi_{\mathrm{sp,l}}^{z}\neq 0$, contributions to the numerator of the bias in the spatial model are multiplied by low weights $w_{i}$, while the denominator of the bias is still relatively large due to the non-spatial information in $\mathbf{x}$. Therefore we expect the bias to be small. This is confirmed in the results shown in Figure 1 (top right) where the bias in the spatial model remains close to zero, even when $\xi_{\mathrm{sp,l}}^{z}$ is increased. In contrast, swapping frequencies makes no difference to the non-spatial model as all frequencies have equal weight and, indeed, the behaviour for this model is the same as in Scenario 1. Web Appendix 3.1 confirms, as before, that the spatial model is superior in terms of fit.
@@ 4.3 Scenario 3: Comparison to the non-spatial model
In Section 3.3 we noted that difference or similarity between the estimates in the spatial and non-spatial models is not directly linked to the existence or otherwise of bias in the spatial model. This is illustrated in Scenarios 1 and 2 above, as in both of these scenarios the estimates in the two models differ but, in Scenario 1, the spatial model is biased whereas, in Scenario 2, it is not. To illustrate this point further, here we simulate two scenarios in which the estimates in the two models are similar, but in one scenario both models are biased whereas in the other they are both largely unbiased. Specifically, in order to generate $\mathbf{z}_{\mathrm{sp,h}}$ and $\mathbf{z}_{\mathrm{sp,l}}$, we fit a reparametrised TPRS spatial effect (see Web Appendix 2) to the Gaussian process with $\kappa=0.1$ where the corresponding 75 highest frequencies are used to generate $\mathbf{z}_{\mathrm{sp,h}}$ and the remaining frequencies to generate $\mathbf{z}_{\mathrm{sp,l}}$. We consider: (1) biased scenario with $(\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x})=(0,4)$ and $(\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z})=(1,1)$; (2) unbiased scenario with $(\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x})=(0,0.1)$ and $(\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z})=(1,0)$. We use $n$ basis functions in the the data analysis model. In the first scenario, $\mathbf{x}$ has only high frequencies which are all confounded, so the spatial model is biased with weights $w_{i}$ close to $1$, making the bias similar to that of the non-spatial model. In the second scenario, there are no overlapping frequencies between $\mathbf{x}$ and $\mathbf{z}$ so both models are unbiased. The results are shown in Figure 1 (bottom, left).
Section 3.3 also showed that while the non-spatial bias tends to exceed the spatial bias, it could also be the other way around. In both Scenarios 1 and 2, the non-spatial bias always stays above the spatial bias. However, in Web Appendix 3.2, we repeat Scenario 1 with a larger proportion of unconfounded low frequency components in $\mathbf{x}$ so that the correlation between $\mathbf{x}$ and $\mathbf{z}$ is low overall, but high at high frequencies. As expected from our analysis, in this latter scenario, the spatial bias exceeds the non-spatial bias.

### v3

@@ 3.1 Bias in the spatial model
Thus, the bias has the intuitive interpretation as the correlation between $\mathbf{x}$ and $\mathbf{z}$ times the size of $\mathbf{z}$ relative to the size of $\mathbf{x}$ in the metric defined by $\bm{\mathbf{\Sigma}}^{-1}$. That is, in this metric, the bias gets large exactly when the covariate and the confounder are very correlated and the confounder is relatively large compared to the covariate. The sign of the bias is the same as the sign of the correlation, i.e. the sign of $\langle\mathbf{x},\mathbf{z}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}$. The proof of the proposition is in Appendix 1.3.
Combining Proposition 1 with the results of Section 2 we get a better understanding of how the bias behaves. Let $\bm{\xi}^{x}=(\bm{\xi}^{xT}_{\mathrm{ns}},\bm{\xi}^{xT}_{\mathrm{sp}})^{T}$ denote the coordinates of $\mathbf{x}$ in the eigenbasis $\mathbf{U}$ such that $\mathbf{x}=\mathbf{U}\bm{\xi}^{x}$, i.e. a non-zero entry in $\bm{\xi}^{x}$ means that the corresponding basis vector is included in $\mathbf{x}$. Similarly, let $\mathbf{z}=\mathbf{U}\bm{\xi}^{z}$. Since $\mathbf{z}$ is spatial, $\bm{\xi}^{z}_{\mathrm{ns}}=\mathbf{0}$. It is now straightforward to prove the following result (see Appendix 1.4. for the proof).
Corollary 1
@@
The expression for the bias is not symmetric in $\mathbf{x}$ and $\mathbf{z}$. In particular, it is not affected by components in $\mathbf{z}$ for which the corresponding component in $\mathbf{x}$ is 0 (i.e. when $\xi_{\mathrm{sp},i}^{z}\neq 0$ but $\xi_{\mathrm{sp},i}^{x}=0$). Thus, any unmeasured spatial variation in the response variable that is independent of the covariate can be ignored for the purposes of the bias. However, such components may affect the overall fit of the model and the estimates of the parameters $\lambda$ and $\sigma^{2}$ which affect the weights $w_{i}$.
@@ 3.2 The role of spatial frequencies
Due to the differences in the weights $w_{i}$, different spatial frequencies affect the bias differently.
Looking at the numerator in Corollary 1, any confounded spatial component of $\mathbf{x}$ (i.e. those with $\xi_{\mathrm{sp},i}^{x}\neq 0$ and $\xi_{\mathrm{sp},i}^{z}\neq 0$ ) contribute to the bias, and the larger the weight, the larger the contribution. So confounding at high frequencies (where weights are close to $1$) induce larger bias, while confounding at low frequencies may give a negligible contribution. Confounding at unpenalised spatial frequencies does not lead to bias (as they have weights 00), and non-spatial components in $\mathbf{x}$ do not contribute to the numerator either. For all confounded frequencies in $\mathbf{x}$ with $w_{i}\neq 0$, the contribution to the bias is proportional to the size $\xi_{\mathrm{sp},i}^{z}$ of the confounder at that frequency. Contributions to the bias from different eigenvectors can partially cancel each other out as they can have different signs.
@@ 3.3 Comparison to the non-spatial model
Irrespective of the size of bias, the non-spatial model (4) is not an appropriate analysis model for the data generation process (2), as the residual spatial variation in the response data violates the assumptions of the model. Indeed, this is the reason for including spatial random effects. Nevertheless, we consider how the bias in model (4) compares to the spatial model (3) as this has been a focus of spatial confounding research in the past. The following corollary is proved in Appendix 1.8.
Corollary 2
@@
Thus, the bias in the non-spatial model differs to that of the spatial model only in that all spatial frequencies have weights $1$ rather than $w_{i}$. Broadly speaking, bias in the non-spatial model reflects the overall correlation between $\mathbf{x}$ and $\mathbf{z}$, whereas bias in the spatial model depends mostly on the correlation at high frequencies. However, as the weights affect both the numerator and denominator of the bias, in general, the relationship between the two models is not straightforward. A detailed computation of the difference and how it is affected by different types of behaviours is given in Appendix 1.4. Although the bias tends to be larger in the non-spatial model than in the spatial model, there are scenarios where it is the other way around.
In the past, the estimate in model (4) was often mistakenly assumed to be unbiased and, therefore, a large difference between this and the estimate in (3) was seen as a sign of confounding bias in the spatial model (and, conversely, a small difference a sign of no confounding). However, Corollary 2 shows that a difference between these estimates tends to arise when the covariate is dominated by low frequencies (as these have weights most different from $1$) and the existence of such a difference does not necessarily mean that there is confounding bias in the spatial model. Conversely, we see that if the covariate is dominated by high frequencies (with weights close to $1$), the estimates in the two models will be similar, but this does not necessarily mean that they are unbiased.
@@ 3.4 Dependence on the smoothing parameter
As bias in the spatial model is caused by smoothing, intuitively, we would expect larger values of the parameter $\lambda$ to lead to more bias when all other components are held fixed. However, as $\lambda$ affects both the numerator and the denominator of the expression in Corollary 1, there can be situations where this is not the case. In Appendix 1.5 we provide a detailed analysis which shows that, although the behaviour of the bias varies, when there is sufficient non-spatial or unconfounded high frequency components in $\mathbf{x}$, the size of the bias broadly increases with $\lambda$ (in line with intuition). The following corollary is also proved in the Appendix 1.6.
Corollary 3
@@
Thus, in the limit $\lambda\rightarrow\infty$, i.e. an entirely smoothed spatial effect, the bias in the spatial model agrees with that of the non-spatial model, except that any contributions from unpenalised spatial components of $\mathbf{x}$ are 00. In the limit $\lambda\rightarrow 0$, i.e. without smoothing, the behaviour of the bias depends on whether or not the covariate has non-spatial information. If non-spatial information is present, the bias can be eliminated by setting $\lambda=0$, though this is usually undesirable as it is likely lead to an inferior model fit. Similarly, if the proportion of unconfounded high frequency components in $\mathbf{x}$ is large, then the size of the bias reduces as $\lambda\rightarrow 0$ and the limit is close to 00. But in the absence of non-spatial or unconfounded high frequency information, the limit could become large and even exceed the bias of the non-spatial model.
@@ 4 Simulation studies
where $\beta\in\mathbb{R}$, $\bm{\epsilon}^{y}\sim N(\mathbf{0},\sigma^{2}\mathbf{I})$, $\bm{\epsilon}^{x}\sim N(\mathbf{0},\sigma^{2}_{x}\mathbf{I})$ for $\sigma,\sigma_{x}>0$. In this specification, $\mathbf{z}$ and $\mathbf{z}^{x}$ are constructed as linear combinations of two spatial fields $\mathbf{z}_{\mathrm{sp,l}}$ and $\mathbf{z}_{\mathrm{sp,h}}$, which are low and high frequency, respectively, i.e.
$\mathbf{z}=\xi_{\mathrm{sp,l}}^{z}\mathbf{z}_{\mathrm{sp,l}}+\xi_{\mathrm{sp,h}}^{z}\mathbf{z}_{\mathrm{sp,h}},\qquad\mathbf{z}^{x}=\xi_{\mathrm{sp,l}}^{x}\mathbf{z}_{\mathrm{sp,l}}+\xi_{\mathrm{sp,h}}^{x}\mathbf{z}_{\mathrm{sp,h}}$
for $\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z},\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x}\in\mathbb{R}$, while $\bm{\epsilon}^{x}$ is spatially unstructured and represents the non-spatial component of $\mathbf{x}$. Unless stated otherwise, we let $\beta=0.5$, $\sigma_{x}=\sigma=1$.
For the regressions in this section, we represent the spatial effect as a thin plate regression spline (TPRS), a computationally efficient reduced rank version of a thin plate spline (Wood,, 2003). The data generating process matches the analysing spatial model which is therefore correctly specified. In Appendix 4, we repeat the simulation where the data generating process is a Gaussian process, so in addition to the spatial confounding bias, there may be further mis-specification bias. In order to generate $\mathbf{z}_{\mathrm{sp,l}}$ and $\mathbf{z}_{\mathrm{sp,h}}$, we start by considering a mean-zero Gaussian process $\bm{\gamma}$ with exponential covariance structure following $C(h)=\exp(-h/\kappa)$ such data $h=\lVert\mathbf{s}-\mathbf{s}^{\prime}\rVert$ for $\mathbf{s},\mathbf{s}^{\prime}\in[0,1]\times[0,1]$, which is assumed for simplicity to have variance of 1. We set $\kappa=0.1$ corresponding to a spatial range of approximately 0.3. We fit thin plate regression splines to the generated $\bm{\gamma}$ with 10 and 800 basis functions, respectively, and use the corresponding fitted values as $\mathbf{z}_{\mathrm{sp,l}}$ and $\mathbf{z}_{\mathrm{sp,h}}$. Thin plate regression spline models are implemented in the R-package mgcv. We use this implementation with generalised cross-validation as the smoothness selection criterion to compare the results of models fitted to simulated data for which we know the true underlying covariate and spatial dependence. We consider 600 basis functions for the thin plate regression splines in the data analysing model.
@@ 4.1 Scenario 1: Confounding at high frequencies
As expected, Figure 1 (top left) shows significant bias in both models under this scenario, increasing linearly in size with $\xi_{\mathrm{sp,h}}^{z}$. The mean squared error (MSE) of fitted values compared to the true expectation of $\mathbf{y}$ is shown in Appendix 3.1 and is always lower for the spatial model, indicating a better fit, which is not surprising as the spatial model is correctly specified, whereas the non-spatial model is mis-specified.
Figure 1: Bias of $\widehat{\beta}$ and $\widehat{\beta}_{\mathrm{ns}}$ in the spatial and non-spatial (ns) models under Scenario 1 (top, left), Scenario 2 (top, right), Scenario 3 (bottom, left), Scenario 4 (bottom, right). The true value of $\beta$ is 0.5.
@@ 4.2 Scenario 2: Confounding at low frequencies
We reconsider Scenario 1 but with high and low frequencies swapped, i.e. $(\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x})=(1,1)$, $\xi_{\mathrm{sp,h}}^{z}=0$ and $\xi_{\mathrm{sp,l}}^{z}\in\{0.2,0.5,1\}$. Since confounding is now at low frequencies, i.e. $\xi_{\mathrm{sp,l}}^{x}\xi_{\mathrm{sp,l}}^{z}\neq 0$, contributions to the numerator of the bias in the spatial model are multiplied by low weights $w_{i}$, while the denominator of the bias is still relatively large due to the non-spatial information in $\mathbf{x}$. Therefore we expect the bias to be small. This is confirmed in the results shown in Figure 1 (top right) where the bias in the spatial model remains close to zero, even when $\xi_{\mathrm{sp,l}}^{z}$ is increased. In contrast, swapping frequencies makes no difference to the non-spatial model as all frequencies have equal weight and, indeed, the behaviour for this model is the same as in Scenario 1. Appendix 3.1 confirms, as before, that the spatial model is superior in terms of fit.
@@ 4.3 Scenario 3: Comparison to the non-spatial model
In Section 3.3 we noted that difference or similarity between the estimates in the spatial and non-spatial models cannot in itself be used to diagnose the existence or otherwise of bias in the spatial model. This is illustrated in Scenarios 1 and 2 above, as in both of these scenarios the estimates in the two models differ but, in Scenario 1, the spatial model is biased whereas, in Scenario 2, it is not. To illustrate this point further, here we simulate two scenarios in which the estimates in the two models are similar, but in one scenario both models are biased whereas in the other they are both largely unbiased. Specifically, in order to generate $\mathbf{z}_{\mathrm{sp,h}}$ and $\mathbf{z}_{\mathrm{sp,l}}$, we fit a reparametrised TPRS spatial effect (see Appendix 2) to the Gaussian process with $\kappa=0.1$ where the corresponding 75 highest frequencies are used to generate $\mathbf{z}_{\mathrm{sp,h}}$ and the remaining frequencies to generate $\mathbf{z}_{\mathrm{sp,l}}$. We consider: (1) biased scenario with $(\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x})=(0,4)$ and $(\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z})=(1,1)$; (2) unbiased scenario with $(\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x})=(0,0.1)$ and $(\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z})=(1,0)$. We use $n$ basis functions in the the data analysis model. In the first scenario, $\mathbf{x}$ has only high frequencies which are all confounded, so the spatial model is biased with weights $w_{i}$ close to $1$, making the bias similar to that of the non-spatial model. In the second scenario, there are no overlapping frequencies between $\mathbf{x}$ and $\mathbf{z}$ so both models are unbiased. The results are shown in Figure 1 (bottom, left).
Section 3.3 also showed that while the bias in the non-spatial model tends to exceed that of the spatial model, it could also be the other way around. In both Scenarios 1 and 2, the bias in the non-spatial model always stays above the spatial model bias. However, in Appendix 3.2, we repeat Scenario 1 with a larger proportion of unconfounded low frequency components in $\mathbf{x}$ so that the correlation between $\mathbf{x}$ and $\mathbf{z}$ is low overall, but high at high frequencies. As expected from our analysis, in this latter scenario, the bias in the spatial model exceeds that of the non-spatial model.